OpenStudy (wade123):

calc help please! (:

OpenStudy (wade123):

can you please help me solve it step by step ? @Jhannybean

OpenStudy (anonymous):

Well do you notice the d/dx infront of the integral?

OpenStudy (wade123):

yess

OpenStudy (anonymous):

I'm not sure if you've learned this yet or not, but this is an example of the fundamental theorem of calculus.

OpenStudy (anonymous):

if you integrate something and then differentiate it what would you get?

OpenStudy (anonymous):

\[\frac{d}{dx}\int\limits_{a}^{b}f(x)=F(b)-F(a)\]

OpenStudy (anonymous):

sorry should be lower case f

OpenStudy (anonymous):

do you follow me?

OpenStudy (wade123):

yess sorry i was looking over it

OpenStudy (anonymous):

so by the fundamental theorem of calculus what would say this should evaluate to?

OpenStudy (wade123):

f(x^3)-f(2)

OpenStudy (anonymous):

yes

OpenStudy (wade123):

:)

OpenStudy (anonymous):

did you get the correct answer?

OpenStudy (wade123):

oh thats it? pretty easy

OpenStudy (anonymous):

well I believe the question is designed to get you to think about what is actually happening here

OpenStudy (anonymous):

namely, the integral is the anti derivative

OpenStudy (jhannybean):

\[\frac{d}{dx}\int\limits_2^{x^2}\ln(x^2)dx=\frac{d}{dx}\left(\ln(x^2)dx\right)|_{2}^{x^2}\]

OpenStudy (jhannybean):

That's how I would write it out.

OpenStudy (jhannybean):

Whoop, without that dx there.

OpenStudy (wade123):

but it would come out to the correct answer right??

OpenStudy (anonymous):

I think the point of this problem is to show that no computation is required if the underlying concept is understood

OpenStudy (anonymous):

yes you would

OpenStudy (wade123):

got it(: thank you guys!(:

OpenStudy (jhannybean):

so \[d(\ln(x^2)) = \frac{1}{x^2} \cdot 2x = \frac{2}{x}\]\[=\left.\frac{2}{x}\right|_{2}^{x^2}\]

OpenStudy (anonymous):

@Jhannybean I think there might be an error in your first statement. It doesn't seem like you actually evaluated the integral.

OpenStudy (anonymous):

maybe I'm missing something

OpenStudy (anonymous):

anyways, do you understand the concept, wade?

OpenStudy (wade123):

yes!

OpenStudy (anonymous):

okay, great!